ZEKE-PFI Spectroscopy of the Sodium-Water and Sodium-Ammonia Dimers

\underline{Geofferey A. Blake}, David A. Rodham, Division of Geological and Planetary Sciences California Institute of Technology MS 170-25 Pasadena, CA 91125.

Single-photon ZEKE-PFI experiments have been performed on the Na-H_{2}O, Na-D_{2}O, Na-NH_{3}, and Na-ND_{3} clusters in a molecular beam, and two-photon ZEKE-PFI experiments have been performed on the Na-NH_{3} and Na-ND_{3} clusters. From the single photon spectra the ionization potentials and intermolecular stretching vibrational frequencies of the cluster ions have been determined, as well as the changes in the intermolecular bending vibrational frequencies upon ionization. The intermolecular bending vibrational frequencies of the Na-NH_{3} and Na-ND_{3} cluster ions were obtained from the two-photon ZEKE-PFI spectra. Partially resolved rotational structure in the spectra will be discussed.